OpenStudy (imamermaid29):

Malachi took cans to the recycling plant and put them into the can-crusher one at a time. After he had already crushed some of the cans, he started recording the number of cans he crushed per minute. The table shows how many cans Malachi put in the can-crusher at different times. What is the equation that represents Malachi’s situation? Record any fractional answers in decimal form. (y – 65) = (x – ) Malachi’s Situation Minutes 3 5 7 9 Cans crushed 49 65 81 97

OpenStudy (imamermaid29):

@word2

OpenStudy (imamermaid29):

@green_1

OpenStudy (green_1):

@mathmate

OpenStudy (green_1):

@AnimeWover

OpenStudy (mathmate):

Hints The standard linear equation in slope-intercept form is y=mx+c where slope=m=(y2-y1)/(x2-x1) c=the y-intercept, i.e. the value of y when x=0. In this case, we do not directly have the value of y when x=0. Using the value of m found above, we can use the point-slope form of the linear equation: y-y1=m(x-x1) where (x1,y1) is any point on the line. In the present case, it can be obtained from the given table, for example, (x1,y1)=(5,65).